# Grounded 3D-LLM with Referent Tokens

### v1

@@ 3.3 Grounded Language Data
Grounded scene caption data generation. The dataset generation for the Grounded Scene Caption dataset (G-SceneCap) leverages ChatGPT and 2D vision-language models, utilizing dense object annotations from existing 3D scan datasets [16, 66, 7], as illustrated in Fig. 3:
@@
Step 2: Condensing objects in local scenes into a caption. For each enumerated anchor object, we form an initial object set by randomly selecting a group of nearby objects. Their captions and coordinates $(x,y,z)$ are input into GPT-4 for captioning, which requires referencing objects by their IDs in the format “[object_phrase object_ID]” in the caption.
@@
Conversion of existing data. In this work, we employ ChatGPT [1] to extract positive phrases for ScanRefer and Multi3DRef datasets for training and evaluation. We refer to the phrase-grounding datasets as “Grounded ScanRefer” (G-ScanRefer) and “Grounded Multi3DRefer” (G-Multi3DRefer). Additionally, we define “Grounded Detection” (G-Detection) as the transformation of detection labels into a sequence of text prompts, similar to GLIP [45].
Extension to embodied dialogue and planning. Grounded scene captions provide local scene information. Using these, we prompt GPT-4 to generate multi-round dialogues and detailed planning sequences in grounded language format. These dialogues, between a human and an embodied agent, encompass tasks including completing activities, discussing layouts, and conducting object-oriented QA. The data generation prompts are detailed in supplementary files.
Table 2: 3D scene-language dataset comparison. We calculate the ratio of instances to tokens to assess training effectiveness per sentence, using the BERT tokenizer [21] to determine the average token length. For sentences with single-object correspondence, each text is paired with one instance.
Dataset #3D Scan #Language Correspondence Average Tokens Inst/Token
Text-level Object-level
ScanRefer [8] 0.7K 52K Sentence Single 22.6 4.4%
Sr3D [2] 1.5K 84K Sentence Single 13.1 7.7%
Sr3D+ [2] 1.5K 115K Sentence Single 12.9 7.7%
Nr3D [2] 0.7K 42K Sentence Single 14.7 6.8%
Multi3DRefer [83] 0.7K 62K Sentence Multi 19.4 7.2%
ScanScribe [88] 3.0K 278K Sentence Single 66.5 1.5%
Grounded Scene Caption 1.5K 107K Phrase Multi 85.9 11.1%
4 Experiments
4.1 Experimental Setup
Datasets and benchmarks. Our evaluation is conducted on the ScanNet dataset [16], an extensive indoor 3D scene dataset that contains 1,513 multi-modal scenes with 3D point clouds, image sequences, and rich point-level instance segmentation annotations. The dataset is divided into 1,201 scenes for training and 312 scenes for validation, with all subsequent benchmarks adhering to these training/validation splits. The CLASP pre-training involves all grounded language data (Sec. 3.3), including G-SceneCap, G-ScanRefer, G-Multi3DRef, and G-Detection. We conduct our evaluation across several benchmarks: ScanRefer [8] and Multi3DRef [83] for single- and multi-object grounding, ScanNet-200 [60] for object detection, and ScanQA [4] and Scan2Cap [14] for 3D QA and dense captioning tasks. These datasets are converted to instruction formats for instruction-tuning and LLM evaluation. Detailed evaluation metrics are provided in the supplementary files.
Baseline methods. Existing methods for downstream tasks are split into discriminative and generative methods in Table 3. (i) Discriminative models perform closed-set classification via detected object proposals, covering single-object [12, 34, 84] and multi-object grounding [83]. 3D-VisTA [88] pre-trains a unified 3D backbone and conducts task-specific fine-tuning. The top 3D close-set detector, Mask3D [62] is shown for reference. (ii) Generative models cover recent 3D LMMs [72, 33, 11, 31], mainly necessitating task-specific fine-tuning for language grounding, termed Specialist LLMs. LL3DA [11] acts as the generalist, supporting language tasks but limited in localization.
4.2 Main Results
@@
Table 3: Evaluation of 3D scene-level LMMs as a generalist. Entries in gray denote models specialized for specific datasets. The term “Generative Model” denotes generative models with open-ended outputs, unlike discriminative models that focus on closed-set classification. “Zero-shot” denotes models used directly without fine-tuning. “Specialist” and “Generalist” categorize models fine-tuned for particular tasks versus those trained jointly.
@@
Method Generative Model ScanRefer Grd. Multi3DRef Grd. Scan2Cap ScanQA ScanNet-200 Det.
@@
Zero-shot LLMs
@@
LL3DA [11] 36.8 65.2 13.5 76.8
Generalist LLMs
LL3DA [11] 36.0 62.9 13.3 75.7
Grounded 3D-LLM 47.9 44.1 45.2 40.6 35.5 70.6 13.4 72.7 15.1 20.8 23.5
Table 4: Comparison of CLASP with previous referring and detection methods. “Point-supervision” involves using point-level instance segmentation supervision. “End-to-end models” are those that can be trained either fully end-to-end or using separate detectors.
@@
2×2$\times$2 × schedule 54.3 47.4 54.6 49.6 27.5 37.0 41.4
Grounded 3D-LLM. As shown in Table 3, Grounded 3D-LLM, without task-specific fine-tuning, outperforms previous LLM-based models in most metrics, positioning it as a promising candidate for a unified framework in 3D scene understanding. Compared to previous generative models, without task-specific fine-tuning, it achieves an improvement of over 10 points in both single and multiple object grounding. Additionally, it attains comparable outcomes in language-based tasks, achieving the highest CIDEr score of 70.6 in Scan2Cap and matching the top BLEU-4 score of 13.4 in ScanQA, surpassing previous method LL3DA, which is limited to language output without referencing capabilities. Notably, it is the first 3D generative model capable of 3D object detection, distinguishing it from its predecessors.
CLASP. As shown in Table 4, CLASP excels in 3D grounding and detection benchmarks, especially in ScanRefer grounding, Multi3DRef multi-object grounding, and ScanNet-200 detection, highlighting its enhanced phrase grounding capabilities compared to earlier models. Notably, in the Multi3DRef grounding task, CLASP outperforms specialized models by over 8 points, demonstrating its strong multi-object handling. With a doubled training schedule, CLASP exceeds all rivals in grounding accuracy and matches the Mask3D model in detection performance. Note that in this study, we primarily use the 1×1$\times$1 × model for Grounded 3D-LLM, as the 2×2$\times$2 × model, while better at grounding, falls short in subsequent language tasks.
4.3 Ablation Studies
Table 5: Ablation studies on the grounding ability of Grounded 3D-LLM. The term “all data” encompasses the G-ScanRefer, G-Multi3DRef, and G-SceneCap datasets used for pre-training.
# CLASP Pre-training data Output Referent type ScanRefer Grd. Multi3DRef Grd.
Acc@0.25 Acc@0.5 F1@0.25 F1@0.5
(a) $1$-to-Many 6.5 5.8 6.4 5.3
(b) G-ScanRefer $1$-to-Many 43.2 36.6 40.2 36.9
(c) G-SceneCap $1$-to-Many 47.6 43.4 44.3 40.1
(d) All data $1$-to-$1$ 46.5 43.3 44.2 40.7
(e) All data $1$-to-Many 47.9 44.1 45.2 40.6
Table 6: Effects of contrastive language scene pre-training for phrase grounding. $\dagger$ represents using ground-truth boxes.
@@ 4.3 Ablation Studies
Ablation studies on referent token. As shown in Table 5, comparing line (a) with others reveals that directly adapting existing closed-set detectors is ineffective for integrating referent tokens in LLMs. In contrast, CLASP pre-training enhances the interpretation of scene referents in language models. Additionally, we analyze two types of referent tokens: one-to-one (d), where each object is matched with a unique token through Hungarian matching, and one-to-many (e), where a single token represents multiple objects. The latter approach generally yields better results.
@@
Dataset ablations for CLASP. Table 6 directly shows the phrase grounding capabilities of the CLASP model using different data across two key points: (i) Zero-shot text evaluation. Without training on ScanRefer or Multi3DRef language data, our CLASP model (d) proves effective compared to previous models (a) and (b). Notably, despite 3D-VisTA’s data generation adhering to sentence-level grounding similar to ScanRefer, our model exceeds it by 4.9 points in Acc@0.25. Furthermore, in experiment (d), our direct evaluation on the Multi3DRef dataset demonstrates robust multi-object grounding ability. (ii) Multi-dataset evaluation. We further investigate the effects of joint training with grounded language datasets, progressively integrating G-ScanRefer (e), G-Detection (f), and G-Multi3DRef (g). This integration boosts performance, achieving 53.2 in Acc@0.25 on ScanRefer and 51.5 in F1@0.25 on Multi3DRef. Notably, Comparison between (g) and (h) shows that the G-SceneCap dataset consistently raises performance by +2.1 points in ScanRefer Acc@0.25 and +3.0 points in Multi3DRef F1@0.25, highlighting its effectiveness.
5 Limitations and Conclusion
@@
We explore the potential of large multi-modal 3D models to unify various downstream 3D vision tasks in an open-ended generative manner. By interpreting scene referents as language tokens, Grounded 3D-LLM links scene objects or regions with descriptions. We propose a contrastive scene-text alignment method that uses curated datasets to facilitate special token modeling in language. However, it exhibits limitations in performance compared to the pre-trained CLASP model, indicating a disparity between discriminative and generative approaches. Scaling the scene-text data to bridge this gap is an objective for future work.
Society Impact. The method enhances the grounding capabilities of large multi-modal models in 3D environments, generally benefiting downstream applications in robotics. However, it does not comprehensively address all indoor scenarios and may produce hallucinated language outputs, thereby posing risks to system integrity.
@@
The supplementary materials for CLASP are organized as follows:
@@
    A detailed introduction to datasets and evaluation metrics is provided in Sec. A.
@@
    Details on pre-training and instruction-tuning implementations are outlined in Sec. B.
@@
    Additional ablation studies are included in Sec. C.
@@
    Visualizations of results and failure cases for downstream tasks are presented in Section D.
@@
    Details on grounded language data generation, along with visualizations of dataset statistics and spatial relationships, are illustrated in Sec. E.
@@
    Data generation prompts for the Grounded Scene Caption dataset are discussed in Sec. F.
  7. 7.
    Multi-task instruction templates are described in Sec. G.
  8. 8.
    Extension to embodied dialogue and planning data is shown in Sec. H
Appendix A Datasets and Evaluation Details
ScanRefer [8]: This benchmark evaluates the ability to localize 3D objects using natural language descriptions from the ScanNet dataset. We assess prediction accuracy by calculating the percentage of predictions with an Intersection over Union (IoU) above $0.25$ and $0.5$ compared to ground truth boxes, following the official metrics [8]. For a fair comparison with prior works, Grounded 3D-LLM extracts the object with the highest score according to the similarity scores $S_{ref}$, grounding a single object following previous studies [8, 88].
Multi3DRef [83]: This dataset is proposed to address the challenge of localizing a flexible number of objects in real-world 3D scenes using natural language descriptions on the ScanNet dataset. The evaluation for the Multi3DRef dataset [83] involves computing the F1 score at IoU thresholds of $0.25$ and $0.5$ between multiple predictions and ground-truth objects. The official metric uses the Hungarian algorithm to match predicted optimally and ground truth bounding boxes, based on the negative IoU as the cost function. For evaluation of Multi3DRef, we employ a score threshold of $0.3$ to filter the candidate objects.
ScanNet-200 Detection [60]: This dataset offers a comprehensive evaluation for 3D instance segmentation across $200$ classes. Evaluation metrics include mean average precision (AP) across different overlap thresholds: AP at $25\%$ overlap (AP@0.250.25{0.25}0.25), AP at $50\%$ overlap (AP@$0.5$), and AP averaged over overlaps in the range of $[0.5\text{:}0.95\text{:}0.05]$ (AP). These metrics follow the official ScanNet benchmark guidelines [16].
Scan2Cap [14]: The dense captioning benchmark is based on ScanRefer dataset [8], requiring models jointly perform detection and generate captions for each object. The evaluation of the language generation component includes BLEU-4 and CIDEr metrics, weighted by the Intersection over Union (IoU) score with the ground truth boxes, which must be greater than $0.25$ or $0.5$. Recent works [11, 6, 14, 13] employ different strategies for matching detection results with ground-truth annotations. We compute these language metrics following prior works [11, 13].
ScanQA [4]: This dataset is annotated based on ScanNet dataset for 3D Question Answering (3D-QA) task to evaluate the visual and spatial understand of 3D environments. This dataset facilitates the processing of visual data from 3D scans, enabling models to respond to text-based inquiries regarding these scenes. For evaluation, metrics of BLEU-4 and CIDEr metrics are utilized.
In the ScanRefer, Multi3DRef, and Scan2Cap benchmarks, bounding boxes within the axis-aligned coordinate system are typically utilized to calculate the 3D IoU [24]. For comparison with previous studies, we derive 3D bounding boxes by extracting the minimum and maximum values from the predicted masks. This post-processing is adopted even though our model provides a more precise point-level mask.
Appendix B Implementation Details
The training hyper-parameters are detailed in Table A-7. All models are trained on 8 NVIDIA A100 (80G) GPUs.
Table 7: Training hyper-parameters. The instruction fine-tuning batch size comprises 5 scenes, each containing up to 200 instructions.

### v2

@@ 3.3 Generating Grounded Scene-Text Data for Pre-training
Grounded Scene Caption dataset generation. The dataset generation for the Grounded Scene Caption dataset (G-SceneCap) leverages ChatGPT and 2D vision-language models, utilizing dense object annotations from existing 3D scan datasets [81], as illustrated in Fig. 4:
@@
Step 2: Condensing objects in local scenes into a caption. For each enumerated anchor object, we form an initial object set by randomly selecting a group of nearby objects. Their captions and coordinates $(x,y,z)$ are input into GPT-4 for captioning, which requires referencing object phrases with their object IDs in the format “[object_phrase object_ID]” in the caption.
@@
A specific coarse-to-fine scene caption generation example with phrase-to-region correspondence, including step-by-step intermediate outputs, is illustrated in Fig. 10. The data generation prompts are detailed in the supplementary files.
@@
Grounded conversion of existing scene-text data. In this work, we employ ChatGPT [82] to extract positive phrases for ScanRefer and Multi3DRef datasets for training and evaluation. We refer to the grounded datasets as G-ScanRefer and G-Multi3DRefer. The transformation of all detection labels into a fixed text prompt is called G-Detection as GLIP [30].
Figure 6: Converting grounded scene-text data to instruction-following format. This figure shows an instruction-following conversion example for the multi-object grounding task. Note that the referent correspondence is converted from the phrase-to-region correspondence of grounded scene-text annotation.
Comparsion with scene-text datasets. Llava [35] uses GPT-4 to generate multi-modal data based on 2D boxes. Similarly, 3D-LLM [21] relies only on 3D boxes to create instruction-following data. Other 3D works [53, 54] add cross-view information to capture view-level differences and interactions, improving open-vocabulary detection. ScanScribe [25] creates scene annotations from human scene graphs. Sceneverse [83] proposes generating scene graphs automatically, then rephrasing them into captions. Multi3DRef [23] generates multi-object correspondence using ScanRefer annotations.
Beyond these efforts, we focus on capturing noun relationships within natural language rather than generating object-level or region-level captions. Additionally, global-view spatial relationships are considered through GPT (Step 2) and rule-based programming (Step 3). As a result, our generated grounded scene-text data achieves the highest object density, averaging one correspondence per nine tokens, leading to over 1.02 million correspondences, as shown in Table II. This provides rich phrase-level correspondence in natural text and allows for seamless conversion to referent correspondence in the following subsection.
3.4 Converting to Grounded Instruction-Following Format for LLM fine-tuning
In this subsection, we describe the transformation of grounded scene-text data from Sec. 3.3 to the instruction-following format.
Converting 3D vision-language tasks to instruction-following formats. For existing 3D vision-language tasks, we can transform them into instruction-following formats. An example of conversion for multi-object grounding tasks including referent correspondence generation is shown in Fig. 6. These convertible 3D vision-language tasks include single and multi-object grounding, object detection, dense captioning, 3D QA, etc.. For each task, we utilize approximately 10-20 structured task-specific instruction-following templates. Please refer to the supplementary file for details on the task-specific instruction-following templates.
Extension to embodied planning and dialogue. To illustrate the potential applications of referent tokens in daily life and enable multi-round conversations, we prompt GPT-4 to generate embodied dialogue and planning based on the Grounded Scene Caption dataset, rather than using static templates. Unlike 3D-LLM [21], our approach produces not only plain dialogue text but also grounded instruction-following text with interleaved phrase-to-region correspondence (referent tokens). These dialogues adhere to several rules: they occur between a human and an embodied agent and involve tasks such as completing daily activities, discussing layouts, and conducting object-oriented QA based on the given caption. The embodied planning text must generate daily instructions and a step-by-step plan for interacting with objects within the scene caption. Refer to the supplementary files for the prompts used in specific dialogue data generation and the visualized results from our jointly trained model.
@@ 3.4 Converting to Grounded Instruction-Following Format for LLM fine-tuning
TABLE III: Evaluation of 3D scene-level LMMs as a generalist. Entries in gray denote models specialized for specific datasets. “Zero-shot” denotes models used directly without fine-tuning. “Specialist” and “Generalist” categorize models fine-tuned for particular tasks versus those trained jointly. Entries in grey denote the incorporation of ground-truth objects into the model inputs.
@@
Method Generative Model ScanRefer Grd. Multi3DRef Grd. ScanQA Scan2Cap ScanNet-200 Det.
@@
Zero-shot LMMs
@@
LL3DA [20] 13.5 76.8 36.8 65.2
Generalist LMMs
LEO [75] 13.2 101.4 38.2 72.4
LL3DA [20] 13.3 75.7 36.0 62.9
Grounded 3D-LLM 48.6 44.0 44.7 40.8 13.2 75.9 35.0 70.2 12.1 16.8 18.9
4 Experiments
4.1 Experimental Setup
Datasets. We evaluate our model on the ScanRefer [5] (single-object grounding), Multi3DRef [23] (multi-object grounding), ScanNet-200 Detection [85] (object detection/instance segmentation), Scan2Cap [14] (dense object captioning), and ScanQA [10] (3D QA) benchmarks. These benchmarks are built on the ScanNet dataset [81], a large-scale indoor 3D scene dataset with 1,513 multi-modal scenes containing 3D point clouds, image sequences, and point-level instance segmentation annotations. The dataset is split into 1,201 scenes for training and 312 for validation, and all benchmarks follow these splits.
Metrics. We follow prior works and adopt the commonly used metrics for these benchmarks. For ScanRefer and Multi3DRef, we compute the Intersection over Union (IoU) at thresholds of $0.25$ and $0.5$. For ScanNet-200, we report 3D instance segmentation using average precision at $25\%$, $50\%$ overlap, and mean average precision (mAP) with IoU thresholds in the $[0.5\text{:}0.95\text{:}0.05]$ range across 200 classes. For Scan2Cap and ScanQA, we evaluate language generation with BLEU-4 and CIDEr, weighted by IoU above $0.25$ or $0.5$.
To maintain consistency with prior evaluations across the ScanRefer, Multi3DRef, and Scan2Cap benchmarks, we extract 3D bounding boxes by taking the minimum and maximum values from the predicted masks, despite our model producing point-level predictions. For Multi3DRef, we apply a score threshold of $0.3$ to filter the predicted objects.
TABLE IV: Effects of diverse instruction-following templates on language tasks.
Diverse Templates ScanQA Scan2Cap
B-4 C B-4@0.25 C@0.25 B-4@0.5 C@0.5
11.2 70.8 31.5 65.8 29.2 62.1
13.2 75.9 36.9 74.1 35.0 70.2
Implementation Details. Unless otherwise specified, CLASP pre-training is conducted on Grounded Scene Caption, ScanNet-200 detection data, ScanRefer, and Multi3DRef data. 150 object queries via farthest point sampling (FPS) is initialized for CLASP and initialized with pre-trained 3D detector Mask3D [4] to accelerate the training. We incorporate the same per-point mask classification loss $\mathcal{L}_{mask}$ with both sigmoid focal loss and Dice loss for segmentation. The overall loss function for CLASP is defined as:
where $T_{mask}$ is computed using Hungarian matching [4] and $T_{text}$ are the ground-truth mappings computed according to ground-truth phrase-to-region correspondence. $\mathcal{L}_{cls}$ represents the sigmoid focal loss.
During instruction fine-tuning, Grounded 3D-LLM adopts Tiny-Vicuna-1B as the LLM by default, keeping it frozen except for the intermediate projection and LoRA [86] layers. Its vision-language model is initialized with a pre-trained CLASP model. To minimize token costs, only 100 object queries are used. The fine-tuning leverages grounded instruction-following datasets, including ScanRefer, Multi3DRef, ScanNet-200 detection, Scan2Cap, ScanQA, scene captioning, embodied dialogue and embodied planning. An object query is considered a positive match only if its mask IoU with the ground-truth object exceeds $0.3$ otherwise the object will be ignored. The overall language loss is used as follows:
$\mathcal{L}_{LLM}=\mathcal{L}_{\text{lang}}+\mathcal{L}_{ref}(S_{ref},T_{ref}),$ (5)
where $L_{\text{lang}}$ is cross-entropy loss as LLAMA [87] and $L_{ref}$ computes the sigmoid focal loss between $S_{ref}$ and the ground-truth referent correspondence $T_{ref}$ from the grounded annotation.
TABLE V: Grounding ability of Grounded 3D-LLM. The term “all data” encompasses the G-ScanRefer, G-Multi3DRef, and G-SceneCap datasets used for pre-training.
# CLASP Pre-training data <ref> type ScanRefer Grd. Multi3DRef Grd.
Acc@0.25 Acc@0.5 F1@0.25 F1@0.5
(a) One-to-Many 6.5 5.8 6.4 5.3
(b) G-ScanRefer One-to-Many 43.2 36.6 40.2 36.9
(c) G-SceneCap One-to-Many 47.6 43.4 44.3 40.1
(d) All data One-to-One 46.5 43.3 44.2 40.7
(e) All data One-to-Many 47.9 44.1 45.2 40.6
TABLE VI: Comparison of CLASP in phrase grounding tasks. “Point-supervision” involves using point-level instance segmentation supervision. “End-to-end models” are those that can be trained either fully end-to-end or using separate detectors.
@@ 4.1 Experimental Setup
TABLE VII: Effects of pre-training datasets for CLASP. $\dagger$ represents using ground-truth boxes.
@@
4.2 Main Results
Baselines. Existing baseline methods for downstream tasks are categorized into discriminative and generative approaches in Table III.
  • Discriminative models focus on single-task outputs with task-specific heads, covering single-object [9, 7, 6] and multi-object grounding [23]. 3D-VisTA [25] pre-trains a unified 3D backbone followed by task-specific fine-tuning, while Mask3D [4], a SOTA transformer-based 3D detector, is trained with instance segmentation heads.
  • Generative models include recent 3D LMMs [19, 26, 20, 21] without task heads, where specialist models require task-specific fine-tuning, and generalist 3D LMMs [20, 75] are trained in unified language format.
Result analysis. As shown in Table III, Grounded 3D-LLM, without task-specific fine-tuning, outperforms previous LLM-based models in most metrics, positioning it as a promising candidate for a unified framework in 3D scene understanding. Compared to prior 3D LMMs, it achieves the best grounding results, improving by over 10 points in both single and multiple object grounding. Additionally, it attains comparable outcomes in language-based tasks, even against LL3DA, which is limited to text output. Notably, it achieves leading performance with a CIDEr score of 75.9 and a BLEU-4 score of 13.2 in ScanQA. Existing generalist 3D LMMs, such as LEO [75] and LL3DA [20], support language tasks but are limited in localization. Furthermore, leveraging the one-to-many correspondence ability of the referent token, we report that Grounded 3D-LLM performs 3D instance segmentation, achieving 12.1 mAP. Despite its inferior performance compared to prior expert models like Mask3D, it demonstrates the potential to unify pure 3D vision tasks, with further improvements remaining as future work.
4.3 Ablation Studies
Diversity of task-specific instruction-following templates. Typical computer vision tasks such as language grounding, object detection, and dense captioning require instruction-following templates to better adapt to pre-trained weights of large language models. We assess the impact of diverse instruction-following templates on language tasks in Table IV. Results show that diverse instruction-following templates consistently yield significant improvements in tasks like dense captioning and 3D QA. Notably, in template-based captioning tasks (Scan2Cap), both CIDEr@0.25 and CIDEr@0.5 see improvements exceeding $8$ points. Employing a single instruction-following template also leads to unstable results in repeated experiments if diverse templates are not used. Therefore, the diversity of instruction-following templates is crucial for sustaining language capabilities in visual tasks.
Ablation studies of the referent token for grounding. As shown in Table V, comparing line (a) with others indicates that directly adapting existing closed-set detectors is ineffective for integrating referent tokens in Grounded 3D-LLM. Incorporating a grounded scene-text dataset for pre-training boosts performance to $43.2$ Acc@0.25 and $40.2$ F1@0.25. With only G-SceneCap for CLASP pre-training, Grounded 3D-LLM already achieves comparable results to all pre-training data.
@@ 4.3 Ablation Studies
Additionally, we analyze two types of referent tokens: one-to-one (d), where each object is matched with a unique token via Hungarian matching, and one-to-many (e), where a single token represents multiple objects. The latter generally yields better results.
Direct comparison of CLASP in phrase grounding tasks. CLASP performs phrase-level scene-text alignment for pre-training, allowing for direct comparison with previous grounding and detection methods. As shown in Table VI, CLASP excels in 3D grounding and detection benchmarks, particularly in ScanRefer grounding, Multi3DRef multi-object grounding, and ScanNet-200 detection, showcasing its enhanced phrase grounding capabilities compared to earlier models. Notably, in the Multi3DRef grounding task, CLASP outperforms specialized models by over 8 points, demonstrating its strong multi-object handling.
Dataset ablations for CLASP. Table VII further ablates the effects of datasets for the scene-text alignment:
@@
(i) Zero-shot text evaluation. Without training on ScanRefer or Multi3DRef language data, our CLASP model (d) proves effective compared to previous models (a) and (b). Notably, despite 3D-VisTA’s data generation adhering to sentence-level grounding similar to ScanRefer, our model exceeds it by 4.9 points in Acc@0.25. Furthermore, in experiment (d), our direct evaluation on the Multi3DRef dataset demonstrates robust multi-object grounding ability.
(ii) Multi-dataset evaluation. We further investigate the effects of joint training with grounded scene-text datasets, progressively integrating G-ScanRefer (e), G-Detection (f), and G-Multi3DRef (g). This integration boosts performance, achieving 53.2 in Acc@0.25 on ScanRefer and 51.5 in F1@0.25 on Multi3DRef. Notably, Comparison between (g) and (h) shows that the G-SceneCap dataset consistently raises performance by +2.1 points in ScanRefer Acc@0.25 and +3.0 points in Multi3DRef F1@0.25, highlighting its effectiveness.
TABLE VIII: Comparison of cross-modal interactors (CMI).
Cross- interactor types ScanRefer Scan2Cap
Acc@0.25 Acc@0.5 B-4@0.5 C@0.5
One-way 47.2 44.0 34.9 70.4
Bi-directional. 47.9 44.1 35.5 70.6
Comparison of cross-modal interactors. Table VIII ablates two key multi-modal interaction components—cross-modal interactors of Fig. 3: text-to-object one-way CMI and bidirectional CMI. The results indicate that incorporating bidirectional cross-attention modules in cross-modal interactors enhances grounding performance (with a 0.7 improvement in Acc@0.25 on ScanRefer and a 0.6 improvement in BLEU-4@0.5 on Scan2Cap) and dense captioning metrics. This improvement suggests that deeper interactions between modalities provide greater benefits for CLASP.
5 Conclusion
@@
We explore the potential of large multi-modal 3D models to unify various downstream 3D vision tasks in unified language modeling. By interpreting scene referents as special language tokens, Grounded 3D-LLM connects scene objects or regions with language, offering a natural way to localize noun phrases in the 3D environment. We demonstrate its general effectiveness across various 3D vision tasks; however, it shows limitations in performance compared to previous expert models, highlighting the disparity between specialist and generalist approaches. Scaling scene-text data to bridge this gap is an objective for future work.
@@
The supplementary materials are organized as follows:
@@
    Additional ablation studies are included in Sec. 6.
@@
    Hyper-parameters of training are included in Sec. 7.
@@
    Visualizations of results and failure cases for downstream tasks are presented in Section 8.
@@
    Details on grounded scene-text data generation, visualizations of dataset statistics, and data generation prompts, are illustrated in Sec. 9.
@@
    Multi-task instruction-following templates are described in Sec. 10.
@@
    Extension to embodied dialogue and planning data is shown in Sec. 11
6 Additional Ablation Studies
Effects of LLM model sizes. As shown in Table A-IX, we compare the performance of GPT-2, Tiny-Vicuna-1B, Vicuna-7B, and Vicuna-13B. Smaller models like GPT-2 struggle with soft referent token learning. In contrast, Grounded 3D-LLM models, ranging from 1B to 13B parameters, significantly outperform GPT-2 and show marginal improvements on Scan2Cap, though they experience a slight decline on ScanQA. Notably, even with only the projection layer trainable (without LoRA), Grounded 3D-LLM (Vicuna-7B) achieves comparable performance, suggesting that the input scene embeddings are semantically rich and can be extracted by the LLM without intermediate trainable layers. Furthermore, the lack of improvement in language tasks with larger models indicates the current task scope may not fully challenge larger-scale LLMs.
TABLE IX: Comparison of different LLMs. For GPT-2, we train the overall backbones with a smaller language head. In models without LoRA, only the projection layers are trained.
LLM LoRA ScanQA Scan2Cap Trainable Params
B-4 C B-4@0.25 C@0.25 B-4@0.5 C@0.5
GPT-2 8.42 44.6 28.4 45.2 26.7 41.9 169 MB
Tiny-Vicuna-1B 13.2 75.9 36.9 74.1 35.0 70.2 157 MB
Vicuna-7B 13.3 74.1 37.8 74.3 35.4 70.8 364 MB
Vicuna-7B 12.4 70.9 37.4 74.2 34.7 69.9 330 MB
Vicuna-13B 12.9 73.2 38.5 74.6 35.9 70.4 487 MB
7 Implementation Details
The training hyper-parameters are detailed in Table A-X. All models are trained on 8 NVIDIA A100 (80G) GPUs.
TABLE X: Training hyper-parameters. The instruction fine-tuning batch size comprises 5 scenes, each containing up to 200 instructions.